# Bayesian Optimization in Language Space: An Eval-Efficient AI Self-Improvement Framework

### v2

@@ 4.3 Theory: Best-of-N gradient’s Asymptotic Equivalence to UCB gradient
In this section, we prove that the Best-of-$N$ over textual gradients selects, in probability, the ascent direction of the UCB acquisition with exploration weight $\beta$ scaling as $\Theta(\sqrt{\ln N})$ (Theorem 2). What allows us to do rigorous theoretical analysis is that LLM prompt embeddings are invertible and bi-Lipschitz: LLM’s prompt embeddings are known to be injective, and therefore invertible (nikolaou2025language); LLM embeddings are Lipschitz (tang2025understanding), implying they are bi-Lipschitz. Hence, small textual edits can be represented as small changes in the embedding space and vice versa. This enables us to map changes in prompts to changes in embeddings in Euclidean space, and later bring the theoretical guarantees back in the prompt space.
Intuitive explanation.
@@
Denote $x$ to be an embedding of a prompt. Pick a small radius $\varepsilon>0$ around $x$ and sample $N$ nearby edits $x_{i}=x+\varepsilon u_{i}$ with $u_{i}$ on the unit sphere. Score each point
$Y_{i}=\mu(x_{i})+\xi_{i}\,\sigma(x_{i}),\quad\xi_{i}\text{ i.i.d. sub-Gaussian and independent of }\{u_{i}\}.$ (15)
Let $M_{N}=\max_{i}\xi_{i}$ and $q_{N}=F^{-1}(1-1/N)$. Since $M_{N}\simeq q_{N}$ and $q_{N}=\Theta(\sqrt{\ln N})$, Best-of-$N$ acts like picking the sampled $x_{i}$ with the largest
$\mu(x_{i})+q_{N}\,\sigma(x_{i}).$ (16)
On the small sphere around $x$ the function $A_{\beta}(x^{\prime})=\mu(x^{\prime})+\beta\,\sigma(x^{\prime})$ has a single best direction. Call it $v_{\beta}$. Draw a narrow cone around $v_{\beta}$. Then, with a large enough $N$, at least one $u_{i}$ lands in that cone with high probability. The winner comes from that cone. Set $\beta=q_{N}$. Then Best-of-$N$ behaves like UCB with weight $\beta$: it picks the sampled point most similar to the $A_{\beta}$ best direction.
Rigorous derivation.
@@
Let $\mathcal{M}$ be the set of syntactically valid prompts. Suppose that there exists a text embedding $\mathsf{E}:\mathcal{M}\to\mathbb{R}^{d}$, where $x\;:=\;\mathsf{E}(\Pi)\in\mathbb{R}^{d}$ is the embedding of a prompt $\Pi\in\mathcal{M}$. We first state the assumptions motivated by how language model embeddings behave in practice, and then state the main theorem, Theorem 2.
@@
At each gradient step of T-BoN BO, we sample edits that correspond to $u_{1},\ldots,u_{N}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\rho$ on the embedding space $\subset\mathbb{R}^{d}$, where $\rho$ is assumed to have a density bounded below by a positive constant with respect to surface measure. Assumption 1 implies that small textual edits induce near‑linear moves in $\mathbb{R}^{d}$, consistent with the recent findings that prompt embeddings preserve phrase similarity under small changes (reimers2019sentence; he2025position; tang2025understanding; nikolaou2025language).
@@
For the current prompt $\Pi$ with $x=\mathsf{E}(\Pi)$, constants $C<\infty$ and $\varepsilon_{0}>0$ such that for every unit vector $u\in\mathbb{R}^{d}$ and every $\varepsilon\in(0,\varepsilon_{0}]$ the edit operator $\mathrm{Apply}$ realizes a first–order move:
$\mathsf{E}\!\big(\mathrm{Apply}(\Pi,\varepsilon u)\big)=x+\varepsilon u+r(\varepsilon,u),\qquad\sup_{\|u\|=1}\|r(\varepsilon,u)\|\leq C\,\varepsilon^{2}.$ (17)
Next, the assumption 2 is a standard regularity/technical condition: let’s call a function $f:\mathbb{R}^{d}\to\mathbb{R}$ is $C^{1,1}$ if it is continuously differentiable and its gradient is Lipschitz: $\exists L<\infty$ such that $\|\nabla f(y)-\nabla f(z)\|\leq L\|y-z\|$ for all $y,z$.
@@
Let $\mu(x)$ and $\sigma(x)$ denote the mean and standard deviation of the evaluation variable $Y$ at embedding $x$. We assume $\mu,\sigma$ are $C^{1,1}$ locally. We use $g=\nabla\mu(x)$ and $h=\nabla\sigma(x)$.
Assumption 3 specifies a heteroscedastic sub-Gaussian model for candidate scores (with i.i.d. mean-zero, unit-variance, unbounded-support noise independent of the sampled directions) and an ideal Best-of-$N$ choice oracle.
@@
Let $\{\xi_{i}\}_{i=1}^{N}$ be i.i.d. mean-zero, unit-variance, Gaussian random variables with unbounded support, independent of $\{u_{i}\}$. For candidates $x_{i}:=x+\varepsilon u_{i}$ we observe
$Y_{i}\;=\;\mu(x_{i})\;+\;\sigma(x_{i})\,\xi_{i},$ (18)
and the Best-of-$N$ oracle returns $i^{\star}\in\arg\max_{i\in[N]}Y_{i}$.
@@
$\Delta x^{(N)}\;:=\;\mathsf{E}\!\big(\mathrm{Apply}(\Pi,\varepsilon\widehat{u}_{N})\big)-x\;=\;\varepsilon\widehat{u}_{N}+r(\varepsilon,\widehat{u}_{N}).$ (19)
@@
$\frac{\Delta x^{(N)}}{\|\Delta x^{(N)}\|}\;=\;\widehat{u}_{N}\;\xrightarrow{p}\;\frac{\nabla A_{\beta_{N}}(x)}{\|\nabla A_{\beta_{N}}(x)\|}\;=\;\frac{g+\beta_{N}h}{\|g+\beta_{N}h\|}.$ (20)
@@
$A_{\beta_{N}}(x+\Delta x^{(N)})\;\geq\;A_{\beta_{N}}(x)\;+\;\varepsilon\,\|\nabla A_{\beta_{N}}(x)\|\;-\;o_{p}(\varepsilon).$ (21)
5 TextGrad-Best-of-N Bayesian Optimization (T-BoN BO)
We now describe the self-improving AI framework proposed in this paper: TextGrad-Best-of-N Bayesian Optimization (T-BoN BO). As discussed in Section 1, T-BoN BO is designed to implement parallel gradient-based UCB-BO in the discrete space of natural language (Figure 3). Specifically, T-BoN BO implements the parallel gradient-based UCB-BO that reuses the final gradient point as the initialization for the next iteration. Therefore, each initial point gives rise to a trajectory: a sequence of candidate points obtained by repeatedly taking gradient-ascent steps and evaluations starting from that initialization.
@@ 5 TextGrad-Best-of-N Bayesian Optimization (T-BoN BO)
Figure 4: Visual summary of T-BoN BO (Algorithm 1) with $T$ iterations, $J$ trajectories, and $G$ Best-of-NN gradients per each iteration and trajectory.
Algorithm.
Throughout $T$ number of optimization iterations, T-BoN BO runs $J$ trajectories in parallel, starting from a set of initial prompts, denoted $\{\Pi_{0}^{j}\}_{j=1}^{J}$. At each iteration and a trajectory, we fix the hyperparameters: the number of candidates per gradient step ($N$) and the number of gradient steps per iteration ($G$).
At iteration $t=0$, it starts from evaluating $\{\Pi_{0}^{j}\}_{j=1}^{J}$, i.e., the initial systems $\{c_{0}^{j}=\Phi(\Pi_{0}^{j})\}_{j=1}^{J}$, on $\{\mathcal{D}_{\mathcal{X}}(\Pi_{0}^{j})\}_{j=1}^{J}$ to obtain initial scores $\{s_{0}^{j}\}_{j=1}^{J}$. The history is then initialized as $H_{0}=\bigcup_{j=1}^{J}\{(\Pi_{0}^{j},c_{0}^{j},s_{0}^{j})\}$. The shared reflection is initialized as $R_{0}=\emptyset$.
At each iteration $t$, for each trajectory $j\in\{1,\ldots,J\}$, we set $\Pi_{t,0}^{j}$ as $\Pi_{t-1}^{j}$ and proceed as follows:
@@
    Best-of-N gradients. (Detailed in Section 4.2) For each trajectory $j$, repeat steps 1) – 4) for $g=1,\ldots,G$:
@@
      Textual gradient generation: From the current prompt $\Pi_{t,g-1}^{j}$, query the textual-gradient generation operator $\nabla_{\text{text}}(\Pi_{t,g-1}^{j};R_{t-1},M_{\text{critic}})$ independently $N$ times to obtain textual gradients $\{\delta^{j,(i)}_{t,g}\}_{i=1}^{N}$.
@@
      Candidate prompt generation: Apply each textual gradient $\delta_{t,g}^{j,(i)}$ to $\Pi_{t,g-1}^{j}$ for all $i$ to get $\{\Pi_{t,g}^{j,(i)}\}_{i=1}^{N}$.
@@
      Candidate system outcome generation: Produce the system outcomes $\{c_{t,g}^{j,(i)}:=\Phi(\Pi_{t,g}^{j,(i)})\}_{i=1}^{N}$.
@@
      Best-of-N selection: Find the best of $\{(\Pi_{t,g}^{j,(i)},c_{t,g}^{j,(i)})\}_{i=1}^{N}$ using the critic and the shared reflection $R_{t-1}$, and set the winner as $(\Pi_{t,g}^{j},c_{t,g}^{j})$.
@@
    Evaluation. For each trajectory $j$,
@@
      Set $c_{t}^{j}=c_{t,G}^{j}$ and $\Pi_{t}^{j}=\Pi_{t,G}^{j}$.
@@
      Evaluate $c_{t}^{j}$ to get the score $s_{t}^{j}=\mathrm{Eval}(c_{t}^{j},\mathcal{D}_{\mathcal{X}}(\Pi_{t}^{j}))$, where
      $\mathrm{Eval}(c_{t}^{j},\mathcal{D}_{\mathcal{X}}(\Pi_{t}^{j}))\;:=\;\frac{1}{L}\sum_{l=1}^{L}r\!\left(y\left(c^{j}_{t},x_{t}^{j,(l)}\right)\right),\qquad x_{t}^{j,(l)}\sim\mathcal{D}_{\mathcal{X}}(\Pi^{j}_{t}),$ (22)
@@
      Given $\{(\Pi_{t}^{j},c_{t}^{j},s_{t}^{j})\}_{j=1}^{J}$, append it to the shared history $H_{t-1}$ to form $H_{t}$.
@@
      If $s_{t}^{j}>s_{t-1}^{j}$, keep $(\Pi_{t}^{j},c_{t}^{j})$. Otherwise, revert trajectory $j$ to the previous state: $(\Pi_{t}^{j},c_{t}^{j})=(\Pi_{t-1}^{j},c_{t-1}^{j})$ and set $s_{t}^{j}=s_{t-1}^{j}$.
    For 4), we follow textgrad and adopt a simple acceptance rule that controls progress: if $s_{t}\geq s_{t-1}$, the new candidate is adopted; otherwise, the system rolls back to the previous iteration’s outcome, i.e., $(\Pi_{t-1},c_{t-1},s_{t-1})$. After evaluation of all trajectories, $\{(\Pi^{j}_{t},c^{j}_{t},s^{j}_{t})\}_{j=1}^{J}$ are appended to the history $H_{t-1}$ to form $H_{t}$.
@@
    Shared Meta-reflection.
    At the end of each iteration $t$, we update the global reflection
    $R_{t}\leftarrow\textsc{MetaReflect}(H_{t};M_{\text{critic}})$
@@
    using the aggregated history from all trajectories for use at iteration t+1t{+}1. Here, $\textsc{MetaReflect}(H_{t};M_{\text{critic}})$ compresses the full history $H_{t}$ of prompts, generated system outcomes, and evaluation results into a natural language reflection $R_{t}$. $R_{t}$ consists of a small set of actionable rules. In the ad generation example, those rules may take the form “emphasize social context,” “avoid cluttered backgrounds,” or “prefer warm lighting.” This reflection serves as a global memory; in subsequent iterations, the critic LLM takes this reflection as a part of its context when generating textual gradients. (For the prompt used for meta-reflection, see Figure 12 in Appendix A).
@@
    There are many ways to construct a reflection from past results. In our experiments, following li2025test, we instruct the critic model $M_{\text{critic}}$ to compare the five lowest- and five highest-performing outcomes from previous iterations (or, when fewer than ten are available, all outcomes) [^4] Limiting the number of outcomes in the critic model’s context window helps mitigate context rotting (hong2025context)., identify recurring visual and textual patterns, and distill them into a concise set of actionable rules.

### v3

@@ 5.3 Theory: Best-of-N gradient’s Asymptotic Equivalence to UCB gradient
We now prove that the Best-of-N over textual gradients selects, in probability, the ascent direction of the UCB acquisition with exploration weight $\beta$ scaling as $\Theta(\sqrt{\ln N})$ (Theorem 1). What allows us to do rigorous theoretical analysis is that LLM prompt embeddings are invertible and bi-Lipschitz: LLM’s prompt embeddings are known to be injective, and therefore invertible (nikolaou2025language); LLM embeddings are Lipschitz (tang2025understanding), implying they are bi-Lipschitz. Hence, small textual edits can be represented as small changes in the embedding space and vice versa. Consequently, small textual edits correspond to small perturbations in the embedding space, allowing us to analyze Best-of-N gradient steps in $\mathbb{R}^{d}$ and then translate the theoretical guarantees back to the prompt space.
5.3.1 Intuitive Explanation
@@
Denote $e$ to be an embedding of a prompt. Pick a small radius $\varepsilon>0$ around $e$ and sample $N$ nearby edits $e_{i}=e+\varepsilon u_{i}$ with $u_{i}$ on the unit sphere. Next, denote the critic LLM’s ($M_{critic}$ from $\S$5.1) internal scoring of each edit, $Y_{i}$, as follows:
$Y_{i}=\mu(e_{i})+\xi_{i}\,\sigma(e_{i}),$ (16)
where $\mu$ is the (implicit) posterior mean, $\sigma$ is the (implicit) epistemic uncertainty, and $\xi_{i}$ is an i.i.d. sub-Gaussian noise variable independent of $u_{i}$. We note that $Y_{i}$ is in line with the BO scoring function from $\S$4.
Let $M_{N}=\max_{i}\xi_{i}$ and let $q_{N}=F^{-1}(1-1/N)$ be the (11/N)(1-1/N)-quantile of the noise distribution. For Gaussian noise, classical extreme-value theory implies that $M_{N}$ concentrates around $q_{N}$ and that $q_{N}=\Theta(\sqrt{\ln N})$. In other words, as $N$ grows large, the largest noise term among $\{\xi_{i}\}_{i=1}^{N}$ behaves like $q_{N}$ up to lower-order fluctuations. Consequently, the index $i^{\star}$ selected by Best-of-N approximately maximizes
$\mu(e_{i})+q_{N}\,\sigma(e_{i}).$ (17)
Consequently, for large $N$, selecting the Best-of-N index $i^{\star}$ is asymptotically equivalent to selecting the maximizer of UCB-style scores with exploration weight $q_{N}$.
Now consider a small sphere around $e$ in embedding space and the UCB acquisition $A_{\beta}(e^{\prime})=\mu(e^{\prime})+\beta\,\sigma(e^{\prime})$. On this sphere, there is a single dominant ascent direction for $A_{\beta}$, which we denote by $v_{\beta}$ (the direction in which $A_{\beta}$ increases the fastest to first order). Draw a narrow cone around $v_{\beta}$. For sufficiently large $N$, at least one of the sampled directions $u_{i}$ falls inside this cone with high probability, and the Best-of-N winner $e_{i^{\star}}$ comes from this cone. Thus, among the locally sampled perturbations $e_{i}=e+\varepsilon u_{i}$, the Best-of-N rule selects a candidate whose direction is (with high probability, for large enough $N$) very close to the ascent direction of the UCB acquisition function with $\beta_{N}:=q_{N}$.
5.3.2 Rigorous Derivation
@@
Let $\Pi$ be the set of syntactically valid prompts. Suppose that there exists a text embedding $\mathsf{E}:\Pi\to\mathbb{R}^{d}$, where $e\;:=\;\mathsf{E}(\pi)\in\mathbb{R}^{d}$ is the embedding of a prompt $\pi\in\Pi$. We first state the assumptions motivated by how language model embeddings behave in practice, and then state the main theorem, Theorem 1.
@@
At each gradient step of TextBO, we sample edits that correspond to $u_{1},\ldots,u_{N}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\rho$ on the embedding space $\subset\mathbb{R}^{d}$, where $\rho$ is assumed to have a density bounded below by a positive constant with respect to surface measure. Assumption 1 implies that small textual edits induce near‑linear moves in $\mathbb{R}^{d}$, consistent with the recent findings that prompt embeddings preserve phrase similarity under small changes (reimers2019sentence; he2025position; tang2025understanding; nikolaou2025language).
@@
For the current prompt $\pi$ with $e=\mathsf{E}(\pi)$, constants $C<\infty$ and $\varepsilon_{0}>0$ such that for every unit vector $u\in\mathbb{R}^{d}$ and every $\varepsilon\in(0,\varepsilon_{0}]$ the edit operator $\mathrm{Apply}$ realizes a first–order move:
$\mathsf{E}\!\big(\mathrm{Apply}(\pi,\varepsilon u)\big)=e+\varepsilon u+r(\varepsilon,u),\qquad\sup_{\|u\|=1}\|r(\varepsilon,u)\|\leq C\,\varepsilon^{2}.$ (18)
Next, Assumption 2 is a standard regularity condition: function $f:\mathbb{R}^{d}\to\mathbb{R}$ is of $C^{1,1}$ if it is continuously differentiable and its gradient is Lipschitz: $\exists L<\infty$ such that $\|\nabla f(y)-\nabla f(z)\|\leq L\|y-z\|$ for all $y,z$.
@@
Let $\mu(e)$ and $\sigma(e)$ denote the mean and standard deviation of the critic LLM’s internal score of embedding $e$. We assume $\mu,\sigma$ are $C^{1,1}$ locally. We use $g=\nabla\mu(e)$ and $h=\nabla\sigma(e)$.
Assumption 3 specifies a heteroskedastic sub-Gaussian model for candidate scores (with i.i.d. mean-zero, unit-variance, unbounded-support noise independent of the sampled directions) and an ideal Best-of-N choice oracle.
@@
Let $\{\xi_{i}\}_{i=1}^{N}$ be i.i.d. mean-zero, unit-variance, Gaussian random variables with unbounded support, independent of $\{u_{i}\}$. For candidates $e_{i}:=x+\varepsilon u_{i}$, the critic LLM observes its internal scoring
$Y_{i}\;=\;\mu(e_{i})\;+\;\sigma(e_{i})\,\xi_{i},$ (19)
and the Best-of-N selection oracle (often built using pairwise tournaments in practice; see $\S$5.2) returns corresponding $i^{\star}\in\arg\max_{i\in[N]}Y_{i}$.
@@
$\Delta e^{(N)}\;:=\;\mathsf{E}\!\big(\mathrm{Apply}(\pi,\varepsilon\widehat{u}_{N})\big)-e\;=\;\varepsilon\widehat{u}_{N}+r(\varepsilon,\widehat{u}_{N}).$ (20)
@@
$\frac{\Delta e^{(N)}}{\|\Delta e^{(N)}\|}\;=\;\widehat{u}_{N}\;\xrightarrow{p}\;\frac{\nabla A_{\beta_{N}}(e)}{\|\nabla A_{\beta_{N}}(e)\|}\;=\;\frac{g+\beta_{N}h}{\|g+\beta_{N}h\|}.$ (21)
@@
$A_{\beta_{N}}(e+\Delta e^{(N)})\;\geq\;A_{\beta_{N}}(e)\;+\;\varepsilon\,\|\nabla A_{\beta_{N}}(e)\|\;-\;o_{p}(\varepsilon).$ (22)
Thus, each Best-of-N gradient step asymptotically performs a first-order improvement on an implicit UCB acquisition function, with exploration strength $\beta_{N}$ controlled entirely by $N$. The full proof is given in Web Appendix B.
Theorem 1 provides the central conceptual and technical contribution of this paper: it shows that a procedure that looks like a purely heuristic rule in language space implicitly implements the same optimism-driven search direction as UCB BO in the model’s embedding space. This is novel and important because it removes the main obstacle to bringing UCB BO’s evaluation-efficiency guarantees into prompt optimization: we never construct an explicit surrogate model, never compute $\sigma(\cdot)$, and never differentiate a closed-form acquisition function, yet the effective update direction converges to the UCB gradient direction. Moreover, the result establishes a new use case for the ubiquitous Best-of-N principle as an interpretable exploration mechanism, where $N$ becomes a principled “exploration knob” for the exploration weight $\beta_{N}=\Theta(\sqrt{\ln N})$. In short, Theorem 1 is what turns Best-of-N gradient from an empirical recipe into a principled BO-style optimizer: it justifies viewing our language-space self-improvement loop as parallel gradient-based UCB-BO and is the reason TextBO can inherit BO’s evaluation-efficiency benefits.
6 TextBO Algorithm
We now describe the self-improving AI algorithm proposed in this paper: TextBO. TextBO is designed to implement parallel gradient-based UCB-BO in the space of natural language prompts. Concretely, TextBO adopts the parallel gradient-based UCB-BO scheme that reuses the final gradient point from one iteration as the initialization for the next. Thus, each initial prompt gives rise to a trajectory: a sequence of candidate prompts and outcomes generated by repeatedly taking gradient-ascent steps and then evaluating the resulting system outcome.
The overview of the TextBO algorithm is shown in Figure 3. For each iteration $t\in[1,T]$ and each trajectory $j\in[1,J]$, we evaluate one prompt; to determine the one prompt to evaluate, we take $G$ steps of Best-of-N gradients beforehand, without any evaluation. We present the pseudocode of the TextBO algorithm (Algorithm 2) and provide the line-by-line description of the details of the procedure below.
@@ 6 TextBO Algorithm
Figure 3: A pictorial overview of TextBO (Algorithm 2) with $T$ iterations, $J$ trajectories, and $G$ Best-of-N gradient steps per iteration and trajectory.
Input: System $\Phi$; $\{\pi_{0}^{j}\}_{j=1}^{J}$; critic $M_{\text{critic}}$; candidates per step $N$; gradient steps $G$; iterations $T$; $J$ trajectories Output: Optimized triples $\{(\pi^{j}_{T},c^{j}_{T},s^{j}_{T})\}_{j=1}^{J}$ 1 for $j=1$ to $J$ do 2 $c_{0}^{j}\leftarrow\Phi(\pi_{0}^{j})$ 3 $s_{0}^{j}\leftarrow\text{Eval}(c_{0}^{j},\mathcal{D}_{\mathcal{X}}(\pi_{0}))$ 4 $H_{0}\leftarrow\{(\pi_{0}^{j},c_{0}^{j},s_{0}^{j})\}$, $R\leftarrow\emptyset$ 5  6for $t=1$ to $T$ do 7 for $j=1$ to $J$ do 8    $\pi^{j}_{t,0}\leftarrow\pi^{j}_{t-1}$ 9    for $g=1$ to $G$ do 10       for $i=1$ to $N$ do 11          $\delta_{t,g}^{j,(i)}\leftarrow\nabla_{\text{text}}(\pi_{t,g-1}^{j};R,M_{\text{critic}})$ 12          $\pi_{t,g}^{j,(i)}\leftarrow\operatorname{Apply}(\pi_{t,g-1}^{j},\delta_{t,g}^{j,(i)})$ 13          $c_{t,g}^{j,(i)}\leftarrow\Phi(\pi_{t,g}^{j,(i)})$ 14           15       $(\pi_{t,g}^{j},c_{t,g}^{j})\leftarrow\textsc{Best-of-N}(\{(\pi_{t,g}^{j,(i)},c_{t,g}^{j,(i)})\},R_{t-1},M_{\text{critic}})$ 16        17    $c_{t}^{j}\leftarrow c^{j}_{t,G}$, $\pi_{t}^{j}\leftarrow\pi_{t,G}^{j}$ 18    $s_{t}^{j}\leftarrow\text{Eval}(c_{t}^{j},\mathcal{D}_{\mathcal{X}}(\pi_{t}))$ 19    $H_{t}\leftarrow H_{t-1}\cup\{(\pi_{t}^{j},c_{t}^{j},s_{t}^{j})\}$ 20    if $s_{t}^{j}<s_{t-1}^{j}$ then 21       $\pi_{t}^{j}\leftarrow\pi_{t-1}^{j}$, $c_{t}^{j}\leftarrow c_{t-1}^{j}$, $s_{t}^{j}\leftarrow s_{t-1}^{j}$ 22        23 $R_{t}\leftarrow\textsc{MetaReflect}(H_{t})$ 24  return $\{c_{t}^{j_{t}^{*}}\}_{t=1}^{T}$, where $j_{t}^{*}:=\operatorname{argmax}_{j}s_{t}^{j}$ Algorithm 2 TextBO
Initialization ($t=0$).
At iteration $t=0$, we first evaluate all initial prompts. For each trajectory $j=1,\dots,J$, we first generate the initial system outcome $\Phi(\pi_{0}^{j})$. Next, we evaluate it on the corresponding input distribution $\mathcal{D}_{\mathcal{X}}(\pi_{0}^{j})$ to obtain a scalar score $s_{0}^{j}$. The shared history is then initialized as $H_{0}\;=\;\bigcup_{j=1}^{J}\{(\pi_{0}^{j},\Phi(\pi_{0}^{j}),s_{0}^{j})\}$, and and the shared reflection is initialized as $R_{0}=\emptyset$.
Per-iteration update ($t\geq 1$).
At each optimization iteration $t\in\{1,\dots,T\}$, the algorithm updates all trajectories in parallel. For each trajectory $j\in\{1,\dots,J\}$, we set the starting prompt for this iteration to be the final prompt from the previous iteration, i.e., $\pi_{t,0}^{j}\leftarrow\pi_{t-1}^{j}$. We then perform the following three steps.
@@
    Best-of-N gradients phase. (Detailed in $\S$5.2.) For each trajectory $j$, we take $G$ successive Best-of-N textual-gradient steps, which do not involve any evaluation. That is, for $j=1,\ldots J$, for $g=1,\dots,G$:
@@
      Textual gradient generation. From the current prompt $\pi_{t,g-1}^{j}$, query the textual-gradient operator
      $\delta_{t,g}^{j,(i)}\sim\nabla_{\text{text}}(\pi_{t,g-1}^{j};R_{t-1},M_{\text{critic}})$
      independently $N$ times to obtain textual gradients $\{\delta_{t,g}^{j,(i)}\}_{i=1}^{N}$.
@@
      Candidate prompt generation. Apply each textual gradient to the current prompt $\pi_{t,g-1}^{j}$ to obtain candidate prompts $\{\pi_{t,g}^{j,(i)}\}_{i=1}^{N}$:
      $\pi_{t,g}^{j,(i)}\leftarrow\operatorname{Apply}(\pi_{t,g-1}^{j},\delta_{t,g}^{j,(i)}),\qquad i=1,\dots,N.$
@@
      Candidate system outcome generation. For each candidate prompt $\pi_{t,g}^{j,(i)}$, generate the corresponding system outcome:
      $\Phi(\pi_{t,g}^{j,(i)}),\qquad i=1,\dots,N.$
@@
      Best-of-N selection. Using the critic model and the shared reflection $R_{t-1}$, select the candidate that is predicted to be the most promising by the critic model among $\{(\pi_{t,g}^{j,(i)},\Phi(\pi_{t,g}^{j,(i)})\}_{i=1}^{N}$, and set the winner as $(\pi_{t,g}^{j},\Phi(\pi_{t,g}^{j}))$.
    After $G$ gradient steps, trajectory $j$ has an updated prompt-outcome pair $(\pi_{t,G}^{j},\Phi(\pi_{t,G}^{j}))$ for this iteration.
    Geometrically, as illustrated in Figure 4, at iteration $t$ the prompts $\{\pi_{t-1}^{j}\}_{j=1}^{J}$ (equivalently, the initial points $\{\pi_{t,0}^{j}\}_{j=1}^{J}$) act as starting locations in the implicit UCB landscape over prompt space. For each trajectory $j$, the $G$ Best-of-N gradient steps climb this landscape, producing the updated prompts $\{\pi_{t,G}^{j}\}_{j=1}^{J}$ that approximate local maxima of the UCB acquisition function. [^4] For $\{\pi_{t,G}^{j}\}_{j=1}^{J}$ to be considered local optima in the usual gradient-based BO sense, $G$ should be large enough. These local optima are then passed to the evaluation phase (Phase 2).
    Figure 4: Illustration of how TextBO (Algorithm 2) implements Best-of-N gradients phase at iteration $t$. The prompts identified as local optima in this phase are evaluated in the later evaluation phase.
@@
    Evaluation phase. For each trajectory $j$:
@@
      Set the trajectory’s candidate for evaluation to the final output of the Best-of-N gradient phase:
      $\pi_{t}^{j}:=\pi_{t,G}^{j}$
@@
      Evaluate $c_{t}^{j}$ to obtain the scalar score
      $s_{t}^{j}$ $:=\mathrm{Eval}(\Phi(\pi_{t}^{j}),\mathcal{D}_{\mathcal{X}}(\pi_{t}^{j}))=\frac{1}{L}\sum_{l=1}^{L}r\!\left(y\!\left(\Phi(\pi_{t}^{j}),x_{t}^{j,(l)}\right)\right),\;\;\textrm{where}\;\;x_{t}^{j,(l)}\sim\mathcal{D}_{\mathcal{X}}(\pi_{t}^{j}).$ (23)
      Note that this is the step where the evaluation happens; and in each iteration, we conduct $J$ evaluations – one for each trajectory.
@@
      Append the new triple $(\pi_{t}^{j},\Phi(\pi_{t}^{j}),s_{t}^{j})$ to the shared history:
      $H_{t}\leftarrow H_{t-1}\cup\{(\pi_{t}^{j},\Phi(\pi_{t}^{j}),s_{t}^{j})\}.$
@@
      Acceptance rule. If the new score does not improve, revert this trajectory to its previous state:
      $\text{if }s_{t}^{j}\leq s_{t-1}^{j}\text{ then }(\pi_{t}^{j},\Phi(\pi_{t}^{j}),s_{t}^{j})\leftarrow(\pi_{t-1}^{j},\Phi(\pi_{t-1}^{j}),s_{t-1}^{j}).$
    This acceptance rule follows textgrad: it enforces monotone non-decreasing scores along each trajectory by rolling back any non-improving update.
@@
    Shared meta-reflection update.
    At the end of iteration $t$, after all $J$ trajectories have been evaluated and the history updated, we recompute the global reflection:
    $R_{t}\leftarrow\textsc{MetaReflect}(H_{t};M_{\text{critic}}).$
@@
    The operator $\textsc{MetaReflect}(H_{t};M_{\text{critic}})$ compresses the full history $H_{t}$—prompts, system outcomes, and scores—into a concise natural-language reflection $R_{t}$. The reflection consists of a small set of actionable rules summarizing cross-trajectory patterns. In the ad-generation setting, such rules might take the form “emphasize social context,” “avoid cluttered backgrounds,” or “prefer warm lighting.” This reflection serves as a global memory: in subsequent iterations, the critic LLM conditions on $R_{t}$ when producing textual gradients and performing Best-of-N judgments.
@@
    There are many ways to construct such reflections. In our experiments, following li2025test, we instruct $M_{\text{critic}}$ to compare the five lowest- and five highest-performing outcomes from previous iterations (or all available outcomes if fewer than ten exist) [^5] Limiting the number of outcomes in the critic model’s context window helps mitigate context rotting (hong2025context)., identify recurring visual and textual patterns, and distill these patterns into a small number of generalizable rules.